# Simulator Ensembles for Trustworthy Autonomous Driving Testing

### v1

@@ 8 Related Work
In this section, we outline two categories of related work. The first category targets ADAS testing approaches that use one simulator for test generation. The second category reports studies that employ multiple simulators for testing.
@@ 9 Conclusion and Future Work
In our empirical study, we evaluated our approach on testing a DNN-enabled ADAS trained for lane-keeping. We compared our approach in terms of effectiveness and efficiency in identifying simulator-agnostic/valid failures, to single-simulator based testing as well as state-of-the-art testing approach which employs multiple simulators. The results of the study show that combining evaluation results from multiple simulators during testing outperforms single-simulator testing, identifying nearly 100% valid failures. Compared to existing multi-simulator approaches, MultiSim identifies, on average, more valid failures. In terms of efficiency, MultiSim performs similarly to other approaches. Additionally, the study demonstrates that machine learning can be used within MultiSim to predict disagreements between test outcomes across different simulators, increasing efficiency and reducing the variation in the number of valid failures.
Our future work is to extend our study to other use cases including more complex systems under tests such as full-stack ADAS and to investigate the root causes behind the disagreements of evaluation outcomes.
@@ 10.1 Funding
This research was funded by the Bavarian Ministry of Economic Affairs, Regional Development and Energy. This work was partially supported and funded by the National Science Foundation grant number 2349961.

### v2

@@ 8 Related Work
In this section, we outline two categories of related work. The first category targets ADAS testing approaches that use one simulator for test generation. The second category reports studies using multiple simulators for testing.
@@ 9 Conclusion and Future Work
In our empirical study, we evaluated our approach on testing two DNN-enabled ADAS trained for lane-keeping. We compared our approach in terms of effectiveness and efficiency in identifying simulator-agnostic/valid failures, to single-simulator based testing as well as state-of-the-art testing approach which employs multiple simulators. The results of the study show that combining evaluation results from multiple simulators during testing outperforms single-simulator testing, identifying nearly 100% valid failures. Compared to existing multi-simulator approaches, MultiSim identifies, on average, more valid failures. In terms of efficiency, MultiSim performs similarly to other approaches. Additionally, the study demonstrates that machine learning can be used within MultiSim to predict disagreements between test outcomes across different simulators, increasing efficiency and reducing the variation in the number of valid failures.
Our future work is to extend our study to other use cases including more complex systems under tests such as full-stack ADAS and to investigate the root causes behind the disagreements of evaluation outcomes. Another future direction could consist in the integration of the proposed multi-simulator concept within other existing test generators, which can act as a complementary layer that analyzes simulator-induced flakiness and verifies the transferability of detected faults across environments.
@@ 10.1 Funding
This research was funded by the Bavarian Ministry of Economic Affairs, Regional Development and Energy. Matteo Biagiola is partially supported by Fondo Istituzionale per la Ricerca granted by Università della Svizzera italiana (USI).